OpenStudy (anonymous):

14. A taco stand sells tacos for $3.25 each. The stand’s expenses for the day are $210. Write an inequality to represent the amount of tacos they need to sell per day to make a profit, and then solve the inequality.

OpenStudy (anonymous):

65 tacos need to be sold

OpenStudy (jdoe0001):

so if you were selling tacos, and you had expenses of $210 how much would you need to make to show profit?

OpenStudy (dangerousjesse):

\(3.25x>210\) Solve for x :)

OpenStudy (triciaal):

Profit = Sales - Cost when sales = cost 3.25* x = 210 x = 210/3.25 = no of tacos per day to break even (P = 0) in order for P >0 then x >